# Exploring the Vulnerability of the Content Moderation Guardrail in Large Language Models via Intent Manipulation

### v1

@@ 1 Introduction
Building on this progress, a handful of studies have begun applying intent detection techniques Zhang et al. (2025); Guan et al. (2024) to enhance the robustness of llm content moderation guardrails against malicious prompt injections Jin et al. (2024a). By analyzing the underlying intent of user prompts, these techniques have shown promising effectiveness against content-level jailbreaks Jiang et al. (2024); Li et al. (2024). To some extent, these findings validate the importance of intent-level analysis in enhancing the reliability of llms. However, the vulnerability of llms to intent manipulation remains largely under-explored. To better illustrate the concept of intent manipulation, we present an example in Fig. 1.
@@ 2.1 Preliminary Background
Let =$\mathcal{F}=\{f_{1},f_{2},\dots,f_{n}\}$ be a collection of closed-source large language models (llms). For each $f_{i}\in\mathcal{F}$, there exists a content moderation mechanism $\Gamma_{i}:\mathcal{X}\times\mathcal{Y}\to\{0,1\}$, where $\mathcal{X}$ is the space of inquiries and $\mathcal{Y}$ the space of responses, such that:
$\Gamma_{i}(x,\ y)=\begin{cases}0,&\text{if }y\text{ is rejected (deemed unsafe)};\\ 1,&\text{if }y\text{ is allowed for output}.\end{cases}$
@@ 2.2 Our Proposed Framework
0:  Vectors of paraphrased prompts $\bm{Y}_{par}$, jailbreak indicators $\bm{Y}_{j}$, harmful scores $\bm{S}_{h}$, and the iterations for successful jailbreaks $\bm{T}_{j}$.
@@
We present our framework workflow in Fig. 2 and further introduce details in Algo. 1. Given a list of harmful inquiries $\bm{X}$ and the number of iterations $T$ for jailbreak, we first initialize a set of llm-based agents, an auxiliary agent, $f_{A}$, the victim model, $f_{V}$, and a monitoring agent, $f_{M}$, and corresponding prompt aggregators, $g_{A}$, $g_{V}$, and $g_{M}$, each of which contains prefix prompt, $x_{A}$, $x_{V}$, and $x_{M}$ (line 1). After initialization, we create empty vectors $\bm{Y}_{par},\ \bm{Y}_{j},\ \bm{S}_{h},\ \bm{T}_{j}\leftarrow\varnothing$ to collect results (line 2). During the jailbreak, for the $n$-th inquiry, we conduct an iterative process as follows: (i) In the $t$-th iteration, we first integrate the inquiry $x_{(n)}^{(t)}$ with a prefix $x_{A}$ by the prompt aggregator, $g_{A}$, and further employ an auxiliary agent, $f_{A}$, to paraphrase the integrated prompt (line 6). (ii) After paraphrasing, we jailbreak the victim model, $f_{V}$, using the above-paraphrased prompt, $y_{par(n)}^{(t)}$, which is aggregated with a prefix $x_{V}$ via $g_{V}$ before the jailbreak (line 8). (iii) After obtaining a response, $y_{atk(n)}^{(t)}$, from the victim model, we monitor whether jailbreak is successful or not by a jailbreak indicator $y_{j(n)}^{(t)}\in\{0,1\}$ and compute the harmful score $s_{h(n)}^{(t)}\in[0,5]$ using a monitoring agent, $f_{M}$. Similarly, we wrap the response with a prefix, $x_{M}$, by $g_{M}$ before monitoring (line 10). If the jailbreak fails, i.e., $y_{j(n)}^{(t)}=0$ (line 11), we update the prompts for the next iteration (line 13). Otherwise, we store the results and break the inner $FOR$ loop (line 16-20). In the end, we return the results, $\bm{Y}_{par},\ \bm{Y}_{j},\ \bm{S}_{h},\ \bm{T}_{j}$, for further evaluation (line 24).
@@
We traverse $N$ inquiries during the jailbreak. For every inquiry, we run $T$ iterations to jailbreak the victim model. Given an assumption that we take $\mathcal{O}(1)$ to interact with llm-based agents and update prompts or store the results, the total time complexity is $\mathcal{O}(N\cdot T)$. Besides, for each successful jailbreak, we update the paraphrased prompt, jailbreak indicator, and harmful score. Thus, the total space complexity is $\mathcal{O}(N)$.
@@ 3 Experiment
In this section, we first introduce the experimental settings and further present our results in detail.
@@
DeepSeek V3 64K / 8K 0.8 0.95
DeepSeek R1 64K / 8K 0.8 0.95
@@
Victim models.
We evaluate a diverse set of mainstream llms as victim models. GPT-4o Hurst et al. (2024), OpenAI’s flagship multi-modal model, integrates advanced moderation systems that leverage intent classification and response filtering. OpenAI o1 Jaech et al. (2024), a frontier reasoning model, appears to employ stricter intent-aware defenses, such as Chain-of-Thought-based moderation guardrails, making it one of the most resilient to harmful information. OpenAI o3-mini OpenAI (2025), while lightweight and cost-efficient, retains an effective moderation pipeline. In contrast, Gemini 2.0 Pro Pichai et al. (2024) by Google exhibits high resilience to malicious inputs. Claude 3.7 Sonnet Anthropic (2025) by Anthropic demonstrates particularly strong moderation capabilities, likely due to its constitutional AI principles, which emphasize harm minimization and intent alignment. On the other hand, DeepSeek V3 Liu et al. (2024a) is a general-purpose llm known for its cost-effectiveness, while DeepSeek R1 Guo et al. (2025) is tailored toward reasoning tasks with improved performance in multi-turn dialogues.
@@
We compare our proposed framework with the red-teaming methods below. PAIR Chao et al. (2023) is an automatic framework for red-teaming black-box llms. TAP Mehrotra et al. (2024) leverages tree-of-thought reasoning to iteratively refine prompt candidates, minimizing the number of queries required to discover effective jailbreak prompts. ArtPrompt Jiang et al. (2024) bypasses llm’s guardrails by encoding sensitive content into visually ambiguous ASCII art. CipherChat Yuan et al. (2024) investigates the alignment issues from a cryptographic standpoint. FlipAttack Liu et al. (2024b) exploits the vulnerability by constructing left-side noise. Deepinception Li et al. (2024) exploits llm’s personality by constructing nested virtual scenarios that induce llmsto respond to harmful behaviors. ReNeLLM Ding et al. (2024) is an automated jailbreak pipeline consisting of prompt rewriting and scenario nesting. FuzzLLM Yao et al. (2024) applies fuzzing-inspired techniques to synthesize jailbreak prompts using templates, constraints, and diverse question sets to induce unsafe outputs in victim llms.
@@
To evaluate the vulnerability of llms, in this study, we employ a common metric, average jailbreak success rate ($\bar{Y}_{j}$), which is defined as the proportion of successful attacks over all attempts:
@@
To assess the harmful potential of paraphrased inquiries, we further employ the average harmfulness score of paraphrased inquiries ($\bar{S}_{h}$), which is defined as the mean model-assigned harmfulness across all paraphrased prompts:
$\bar{S}_{h}=\frac{1}{N}\sum_{n=1}^{N}\bm{S}_{h(n)},$
where $\bm{S}_{h(n)}$ denotes the harmfulness score of the $n$-th paraphrased inquiry, as predicted by the monitoring agent. Higher $\bar{S}_{h}$ indicates stronger semantic preservation of harmful intent across paraphrases.
@@
None 36.54 2.38 84.04 1.35 48.75 1.65 82.25 1.07 36.00 2.22 79.00 1.76 65.00 2.83 88.12 1.83
ASS 29.04 2.46 87.12 1.55 42.75 2.02 83.25 1.10 36.00 2.28 82.00 1.49 60.62 3.11 86.88 1.82
MSW 29.04 2.26 88.08 1.45 51.75 1.80 88.75 0.84 24.00 2.29 81.00 1.46 54.37 2.60 82.50 1.95
CES 31.54 2.37 85.19 1.68 52.25 1.82 85.25 0.94 37.00 2.26 81.00 1.56 62.50 2.97 88.75 1.88
Table 3: Evaluation of jailbreak performance across different paraphrasing strategies, with and without fuzzy intent matching. GPT-4o is used as the victim model in all experiments.“ASS”, “MSW”, and “CES” denote the paraphrasing strategies: Alter Sentence Structure, Misspell Sensitive Words, and Change Expression Style, respectively.
@@
Table 4: Evaluation of three types of intent-driven jailbreak methods targeting GPT4o across four public benchmarks. Obfuscation (OBF) denotes that we revise inquiries with fuzzy intent. Structuration (STR) denotes that we structure inquiry into an outline using an auxiliary model. Elaboration (ELA) denotes that we elaborate the outline into a detailed statement on the injected prompts to the victim model.
@@
Table 5: Ablation study on the vulnerability of CoT-based defending paradigms in the reasoning model (o1). Building upon our proposed mechanism “STR+ELA” (evaluated in Table 4), we introduce an enhanced variant, “FSTR+SPIN”, specifically designed for reasoning models. Here, “FSTR” involves generating outlines with fuzzy intent via an auxiliary model, and “SPIN” refers to spinning the malicious intent into seemingly benign content.
Conjecture validations and ablation studies.
Discussion of the imperative sensitivity bias.
@@
Table 6: Generalization assessment of our proposed framework (STR+ELA) when evaluated on various victim models across four public benchmarks.
Generalization test.
To assess the generalizability of our proposed framework (STR+ELA), we deploy several mainstream llms as victim models. As presented in Tab. 6, our framework consistently achieves stable jailbreak performance across the majority of victim models, demonstrating strong generalization capabilities. Notably, we observe that two DeepSeek models exhibit relatively weaker moderation guardrails. We argue that this is due to their primary alignment with Chinese-language use cases, which may lead to less stringent moderation and intent recognition for English prompts. In contrast, models such as Claude 3.7 Sonnet and OpenAI’s o3-mini demonstrate stronger resistance to harmful information in English. This resistance may be attributed to advancements in recent alignment. Overall, these results confirm that our framework maintains a high level of transferability and effectiveness across diverse llms’ architectures and alignment strategies.
@@
Table 7: Comparison of our framework and various competing models against the victim model (GPT-4o) across four public benchmarks, under the defending settings with and without intent analysis (IA). We evaluate the jailbreak performance via $\bar{Y}_{j}$.
@@ 4 Related Work
Existing jailbreak attacks against llms can be broadly categorized into white-box and black-box approaches Yi et al. (2024). White-box attacks, such as GCG Zou et al. (2023) and AutoDAN Zhu et al. , rely on access to llms’ gradients or model parameters to craft adversarial prompts. In contrast, black-box attacks do not require access to model internals and instead manipulate input-output content to bypass safety guardrails Lapid et al. (2023). Several black-box methods have been proposed to systematically evade content moderation guardrails of llmXue et al. (2023); Deng et al. (2023). Chao et al. (2023) introduce an automated framework for generating jailbreak prompts in a black-box setting, while Wei et al. (2023) construct jailbreak templates through carefully designed contextual examples. Jiang et al. (2024) leverage ASCII-based transformations to obfuscate sensitive tokens, whereas Xu et al. (2024) propose a context-aware multi-agent framework, RedAgent, for efficient red-teaming. Most of these methods attempt to bypass the guardrails at a content level by transforming harmful content. Recently, however, a small number of studies have shifted the interest to intent-level reasoning Yu et al. (2023); Zhang et al. (2024b); Pu et al. (2024); Xue et al. (2025). For instance, Yao et al. (2024) generate fuzzed variations of jailbreak prompts to cover the malicious intent. Shang et al. (2024) preliminarily reveal that llms’ guardrails will be degraded when the intent of queries is obfuscated.

### v2

@@ 1 Introduction
Building on this progress, a handful of studies have begun applying intent detection techniques Zhang et al. (2025); Guan et al. (2024) to enhance the robustness of llm content moderation guardrails against malicious prompt injections Jin et al. (2024a). By analyzing the underlying intent of user prompts, these techniques have shown promising effectiveness against content-level jailbreaks Jiang et al. (2024b); Li et al. (2024). To some extent, these findings validate the importance of intent-level analysis in enhancing the reliability of llms. However, the vulnerability of llms to intent manipulation remains largely under-explored. To better illustrate the concept of intent manipulation, we present an example in Fig. 1.
@@ 2.1 Preliminary Background
Let $\mathcal{F}=\{f_{1},f_{2},\dots,f_{n}\}$ be a collection of closed-source large language models (llms). For each $f_{i}\in\mathcal{F}$, there exists a content moderation mechanism $\Gamma_{i}:\mathcal{X}\times\mathcal{Y}\to\{0,1\}$, where $\mathcal{X}$ is the space of inquiries and $\mathcal{Y}$ the space of responses, such that:
[…]
@@ 2.2 Our Proposed Framework
0: Vectors of paraphrased prompts $\bm{Y}_{par}$, jailbreak indicators $\bm{Y}_{j}$, harmfulness scores $\bm{S}_{h}$, and the iterations for successful jailbreaks $\bm{T}_{j}$.
@@
We present our framework workflow in Fig. 2 and further introduce details in Algo. 1. Given a list of harmful inquiries $\bm{X}$ and the number of iterations $T$ for jailbreak, we first initialize a set of llm-based agents, an auxiliary agent, $f_{A}$, the victim model, $f_{V}$, and a monitoring agent, $f_{M}$, and corresponding prompt aggregators, $g_{A}$, $g_{V}$, and $g_{M}$, each of which contains prefix prompt, $x_{A}$, $x_{V}$, and $x_{M}$ (line 1). After initialization, we create empty vectors $\bm{Y}_{par},\ \bm{Y}_{j},\ \bm{S}_{h},\ \bm{T}_{j}\leftarrow\varnothing$ to collect results (line 2). During the jailbreak, for the $n$-th inquiry, we conduct an iterative process as follows: (i) In the $t$-th iteration, we first integrate the inquiry $x_{(n)}^{(t)}$ with a prefix $x_{A}$ by the prompt aggregator, $g_{A}$, and further employ an auxiliary agent, $f_{A}$, to paraphrase the integrated prompt (line 6). (ii) After paraphrasing, we jailbreak the victim model, $f_{V}$, using the above-paraphrased prompt, $y_{par(n)}^{(t)}$, which is aggregated with a prefix $x_{V}$ via $g_{V}$ before the jailbreak (line 8). (iii) After obtaining a response, $y_{atk(n)}^{(t)}$, from the victim model, we monitor whether jailbreak is successful or not by a jailbreak indicator $y_{j(n)}^{(t)}\in\{0,1\}$ and compute the harmfulness score $s_{h(n)}^{(t)}\in[0,5]$ using a monitoring agent, $f_{M}$. Similarly, we wrap the response with a prefix, $x_{M}$, by $g_{M}$ before monitoring (line 10). If the jailbreak fails, i.e., $y_{j(n)}^{(t)}=0$ (line 11), we update the prompts for the next iteration (line 13). Otherwise, we store the results and break the inner $FOR$ loop (line 16-20). In the end, we return the results, $\bm{Y}_{par},\ \bm{Y}_{j},\ \bm{S}_{h},\ \bm{T}_{j}$, for further evaluation (line 24).
@@
We traverse $N$ inquiries during the jailbreak. For every inquiry, we run $T$ iterations to jailbreak the victim model. Given an assumption that we take $\mathcal{O}(1)$ to interact with llm-based agents and update prompts or store the results, the total time complexity is $\mathcal{O}(N\cdot T)$. Besides, for each successful jailbreak, we update the paraphrased prompt, jailbreak indicator, and harmfulness score. Thus, the total space complexity is $\mathcal{O}(N)$.
@@ 3 Experiment
In this section, we first introduce the experimental settings and further present our results in detail. Our code is available at https://github.com/junzhuang-code/IntentPrompt.
@@
DeepSeek V3/R1 64K / 8K 0.8 0.95
Qwen3-14B/32B/235B-A22B 32K / 32K 0.8 0.95
Llama 4 Scout 10M / 8K 0.8 0.95
Mixtral-8x7b 32K / 4K 0.8 0.95
@@
llms.
We evaluate a diverse set of mainstream llms as follows. Among closed-source llms, GPT-4o Hurst et al. (2024), OpenAI’s flagship multimodal model, integrates advanced moderation systems that combine intent classification with response filtering. Building on this foundation, OpenAI o1 Jaech et al. (2024), a frontier reasoning model, appears to adopt even stricter intent-aware defenses—such as Chain-of-Thought-based moderation guardrails—making it one of the most resilient models against harmful content. Complementing these high-end systems, the lightweight and cost-efficient OpenAI o3-mini OpenAI (2025) maintains an effective moderation pipeline. In contrast, Google’s Gemini 2.0 Pro Pichai et al. (2024) also exhibits high resilience to malicious inputs. Finally, Claude 3.7 Sonnet Anthropic (2025) from Anthropic showcases particularly robust moderation behavior, likely driven by its constitutional AI principles focused on harm minimization and intent alignment. For open-source llms, DeepSeek V3 Liu et al. (2024a) stands out for its cost-effectiveness, whereas its sibling DeepSeek R1 Guo et al. (2025) is specifically tuned for reasoning. Moreover, we consider Alibaba’s Qwen3 series Yang et al. (2025), including Qwen3-14B, Qwen3-32B, and the 235B-parameter mixture-of-experts (MoE) model Qwen3-235B-A22B with 22B active parameters. Beyond the Qwen3 series, Meta’s Llama4 Scout Meta (2025)—a 109B-parameter MoE llm with 17B active parameters—excels at multi-modal tasks, while Mistral AI’s Mixtral-8x7B Jiang et al. (2024a), a sparse mixture-of-experts (SMoE) model, further broadens our evaluation landscape.
@@
We compare our proposed framework with the red-teaming methods below. PAIR Chao et al. (2023) is an automatic framework for red-teaming black-box llms. TAP Mehrotra et al. (2024) leverages tree-of-thought reasoning to iteratively refine prompt candidates, minimizing the number of queries required to discover effective jailbreak prompts. ArtPrompt Jiang et al. (2024b) bypasses llm’s guardrails by encoding sensitive content into visually ambiguous ASCII art. CipherChat Yuan et al. (2024) investigates the alignment issues from a cryptographic standpoint. FlipAttack Liu et al. (2024b) exploits the vulnerability by constructing left-side noise. Deepinception Li et al. (2024) exploits llm’s personality by constructing nested virtual scenarios that induce llmsto respond to harmful behaviors. ReNeLLM Ding et al. (2024) is an automated jailbreak pipeline consisting of prompt rewriting and scenario nesting. FuzzLLM Yao et al. (2024) applies fuzzing-inspired techniques to synthesize jailbreak prompts using templates, constraints, and diverse question sets to induce unsafe outputs in victim llms.
@@
To evaluate the vulnerability, we employ a common metric, jailbreak success rate ($\bar{Y}_{j}$), defined as the proportion of successful attacks over all attempts:
@@
To assess the harmfulness of paraphrases that bypass guardrails, we define the harmfulness score ($\bar{S}_{h}$) as the average model-assigned harmfulness over all paraphrases in successful jailbreaks:
$\bar{S}_{h}=\frac{\sum_{n=1}^{N}\mathbb{I}\{\bm{Y}_{j(n)}\text{=}1\}\bm{S}_{h(n)}}{\sum_{n=1}^{N}\mathbb{I}\{\bm{Y}_{j(n)}\text{=}1\}},$
where $\bm{S}_{h(n)}$ denotes the harmfulness score of the $n$-th paraphrased inquiry by the monitoring agent. Higher $\bar{S}_{h}$ indicates stronger semantic preservation of harmful intent among successful paraphrases. Importantly, a lower $\bar{S}_{h}$ doesn’t necessarily imply ineffectiveness of the jailbreak strategy, as it may reflect a shift in semantics that reduces the perceived harmfulness.
@@
[…]
Table 3: Evaluation of jailbreak performance against GPT-4o as the victim model by jailbreak success rate $\bar{Y}_{j}$ (%) and harmfulness score $\bar{S}_{h}$ across different paraphrasing strategies, without and with fuzzy intent matching (highlighted by a gray background). “ASS”, “MSW”, and “CES” denote the paraphrasing strategies: Alter Sentence Structure, Misspell Sensitive Words, and Change Expression Style, respectively.
@@
Table 4: Evaluation of three types of intent-driven jailbreak methods targeting GPT4o across four public benchmarks by jailbreak success rate $\bar{Y}_{j}$ (%), harmfulness score $\bar{S}_{h}$, and iterations for successful jailbreaks $\bar{T}_{j}$. Obfuscation (OBF) denotes that we revise inquiries with fuzzy intent. Structuration (STR) denotes that we structure inquiry into an outline using an auxiliary model. Elaboration (ELA) denotes that we elaborate the outline into a detailed statement on the injected prompts to the victim model.
@@
Table 5: Ablation study on the vulnerability of CoT-based defending paradigms in the reasoning model (o1) evaluted by jailbreak success rate $\bar{Y}_{j}$ (%), harmfulness score $\bar{S}_{h}$, and iterations for successful jailbreaks $\bar{T}_{j}$. Building upon our mechanism “STR+ELA” (evaluated in Table 4), we introduce an enhanced variant, “FSTR+SPIN”, specifically designed for reasoning models. Here, “FSTR” involves generating outlines with fuzzy intent via an auxiliary model, and “SPIN” refers to spinning the malicious intent into seemingly benign content.
@@
Qwen3-235B-A22B 94.42 3.38 1.27 95.00 2.64 1.46 96.00 3.47 1.21 98.12 3.33 1.38
Llama4-Scout 80.96 3.48 1.80 82.00 2.72 2.11 92.00 3.26 1.63 90.62 3.61 1.30
Table 6: Generalization test of our variant (STR+ELA) using various victim models across four public benchmarks evaluated by jailbreak success rate $\bar{Y}_{j}$ (%), harmfulness score $\bar{S}_{h}$, and iterations for successful jailbreaks $\bar{T}_{j}$.
Auxiliary Monitoring $\bar{Y}_{j}$ $\bar{S}_{h}$ $\bar{T}_{j}$ Time
Gemini1.5-Flash Gemini1.5-Flash 90.00 3.69 1.43 6.88
Qwen3-14B Gemini1.5-Flash 95.00 2.57 1.14 14.64
Mixtral-8x7b Gemini1.5-Flash 91.00 3.40 1.54 10.83
Gemini1.5-Flash GPT4o-mini 96.00 4.02 1.24 5.19
Gemini1.5-Flash Qwen3-32B 94.00 3.86 1.19 16.77
Table 7: Generalization test of our variant (STR+ELA) evaluated on various auxiliary and monitoring agents (the victim model is GPT-4o) on the JailbreakBench dataset by metrics, including jailbreak success rate $\bar{Y}_{j}$ (%), harmfulness score $\bar{S}_{h}$, iterations for successful jailbreaks $\bar{T}_{j}$, and average runtime per inquiry (in seconds).
Conjecture validations and ablation studies.
Discussion of the imperative sensitivity bias.
@@
Table 8: Comparison of our framework and various competing models against the victim model (GPT-4o) across four public benchmarks, under the defending settings with and without intent analysis (IA). We evaluate the performance by jailbreak success rate $\bar{Y}_{j}$.
Generalization test.
To assess the generalizability of our proposed framework (STR+ELA), we deploy several mainstream llms that are known for their strong generative capabilities as victim models. As presented in Tab. 6, our framework consistently achieves stable jailbreak performance across the majority of victim models, demonstrating strong generalization capabilities. Notably, we observe that two DeepSeek models exhibit relatively weaker moderation guardrails. We argue that this is due to their primary alignment with Chinese-language use cases, which may lead to less stringent moderation and intent recognition for English prompts. In contrast, models such as Claude 3.7 Sonnet and OpenAI’s o3-mini demonstrate stronger resistance to harmful information in English. This resistance may be attributed to advancements in recent alignment. Furthermore, we assess how varying the auxiliary and monitoring agents affects overall performance. As shown in Tab. 7, replacing Gemini-1.5-Flash (Around 13.6B size) with other llms leads to consistent improvements. This is possible as stronger models contribute to more effective generation and evaluation. We also observe that the average runtime is significantly affected by the API call rate limits of different models. Overall, these results confirm that our framework maintains a high level of transferability and effectiveness across diverse llms’ architectures and alignment strategies.
@@
ArtPrompt CipherChat FlipAttack Deepinception ReNeLLM FuzzLLM Ours
12.17 24.05 11.48 19.80 60.23 23.85 6.88
Table 9: Comparison of average runtime per inquiry (in seconds) between various competing methods and our variant (STR+ELA).
Runtime comparison.
We further compare the average runtime per inquiry (in seconds) between competing methods and our framework (STR+ELA) under the same model settings, indicating a high efficiency in our framework.
$T$ 1 2 3 4 5 6
$\bar{Y}_{j}$ 64.00 84.00 87.00 88.00 91.00 89.00
$\bar{S}_{h}$ 3.33 3.28 3.45 3.48 3.60 3.49
$\bar{T}_{j}$ 1.00 1.14 1.18 1.23 1.31 1.36
Time 4.09 7.55 7.61 8.27 8.10 9.06
Table 10: Effect of the maximum number of iterations $T$ on our variant (STR+ELA) using GPT-4o as the victim model evaluated by jailbreak success rate $\bar{Y}_{j}$ (%), harmfulness score $\bar{S}_{h}$, iterations for successful jailbreaks $T_{j}$, and average runtime per inquiry (in seconds).
Analysis of hyperparameters.
As presented in Tab. 10, we analyze how varying the maximum number of iterations $T$ affects our framework using GPT-4o as the victim model on the JailbreakBench dataset as an example, highlighting the trade-off between iteration budget and performance. Considering higher $\bar{Y}_{j}$ and lower runtime, we choose $T=5$ in this study.
@@ 4 Related Work
Given that llms are widely applied in various domains Zhuang and Guan (2025), existing jailbreak attacks against llms can be broadly categorized into white-box and black-box approaches Yi et al. (2024); Verma et al. (2024). White-box attacks, such as GCG Zou et al. (2023) and AutoDAN Zhu et al. , rely on access to llms’ gradients or model parameters to craft adversarial prompts. In contrast, black-box attacks do not require access to model internals and instead manipulate input-output content to bypass safety guardrails Lapid et al. (2023). Several black-box methods have been proposed to systematically evade content moderation guardrails of llmXue et al. (2023); Deng et al. (2023); Yadav et al. (2025). Chao et al. (2023) introduce an automated framework for generating jailbreak prompts in a black-box setting, while Wei et al. (2023) construct jailbreak templates through carefully designed contextual examples. Jiang et al. (2024b) leverage ASCII-based transformations to obfuscate sensitive tokens, whereas Xu et al. (2024) propose a context-aware multi-agent framework, RedAgent, for efficient red-teaming. Most of these methods attempt to bypass the guardrails at a content level by transforming harmful content. Recently, however, a small number of studies have shifted the interest to intent-level reasoning Yu et al. (2023); Zhang et al. (2024b); Pu et al. (2024); Xue et al. (2025). For instance, Yao et al. (2024) generate fuzzed variations of jailbreak prompts to cover the malicious intent. Shang et al. (2024) preliminarily reveal that llms’ guardrails will be degraded when the intent of queries is obfuscated.